OpenStudy (anonymous):

If 1/2x +1/2(1/2x + 1/2(1/2x +1/2(1/2x + ... = y, then x = ?

OpenStudy (phi):

is the x in the denominator or numerator. Is it \[ \frac{1}{2x} or \frac{1}{2}x ?\]

OpenStudy (phi):

Let's assume it is (1/2)x = \( \frac{1}{2} x \) Notice you can rewrite the sum as \[ y= \frac{1}{2}x + \frac{1}{4}x + \frac{1}{8}x + ... \] or \[ y= x(\frac{1}{2} + \frac{1}{4} + \frac{1}{8} + ... )\] Call the sum in the parentheses S: \[ S=\frac{1}{2} + \frac{1}{4} + \frac{1}{8} + ... \] and \[ \frac{1}{2}S=\frac{1}{4} + \frac{1}{8} + ... \] subrtract the two series to get \[ \frac{1}{2}S= \frac{1}{2} \] or S=1 so we find \[ y=x \]